1) The length of a square's side is 10 cm. This square's perimeter is 7 cm more than the perimeter of an equilateral triangle.
a) Let's represent the length of one side of the triangle. What equation models this situation.
b) Solve the equation to find the length of the triangle's sides. Verify your answer.

Answer by rothauserc(4718)

let s be the length of one side of the equilateral triangle
:
a) 3s + 7 = 40
s = (40 - 7) / 3
:
b) s = 33 / 3 = 11
:
**********************************************************
the length of a side of the equilateral triangle is 11 cm
check answer
:
3 * 11 + 7 = 40
:
33 + 7 = 40
:
40 = 40
:
answer checks
**********************************************************
: